Suppose the population over 15 is 20 million, the labour force Participation rate is 60%, and the unemployment rate is 5%. If the labour force grows by 3% and employment grows by 2%, what does the unemployment rate become?
1
Expert's answer
2019-06-04T12:01:41-0400

To calculate the rate of unemployment in this case, it is important to first get labor force, number of the unemployed, number of the employed before and after the growth.

The labor force is 60% of the population above 15.

So labor force: 60% ×20m= 12 Million

The people who are unemployed: 5% × 12M= 0.6M

Therefore, the number of people employed is 12M – 0.6M = 11.4 M

If the labor force grows by 3%. The increase in labor force is; 3% × 12M=0.36M

The new labor force will be; 12M + 0.36M= 12.36M

If employment grows by 2%. The increase in employment is 2% × 11.4M= 0.228M

The number of people employed will be; 11.4M + 0.228M= 11.628M

People unemployed is; 12.36M- 11.628M = 0.732M

The formula for unemployment rate is;

People unemployed/ Total Labor force × 100%

= 0.732/12.36 × 100

= 5.9%

Therefore, the unemployment rate would become 5.9 %.